# Scalable Spatio-Temporal SE(3) Diffusion for Long-Horizon Protein Dynamics

### v1

@@ 3.2 Architecture
This architectural choice is key to our model’s scalability. We analyze the per-layer computational complexity of our ST attention () against general ST (i.e., “space-then-time”) architectural paradigms (see Section˜8 for details). Most baseline architectures employ a Pairformer [jumper2021_af2] backbone, incurring a cubic spatial complexity due to expensive triangular attention operations. We consider two representative baseline configurations: (1) Pairformer + Pair Temporal Attention (e.g., ConfRover [shen2025simultaneous_confrover]): This configuration scales as . Here, the pairwise temporal attention alone incurs cost, matching that of STAR-MD, but the additional cubic spatial overhead makes it significantly more expensive. Furthermore, this approach incurs a quadratic memory cost for caching pairwise key-value states during autoregressive generation, which becomes a severe bottleneck for long rollouts. In contrast, STAR-MD only requires memory for caching single-residue features. (2) Pairformer + Single Temporal Attention: This configuration scales as . While the temporal scaling is theoretically superior to STAR-MD’s in the limit , the cubic spatial cost dominates in practical regimes. STAR-MD thus offers a favorable trade-off, enabling efficient modeling of long trajectories for large proteins without the cubic spatial overhead.
@@ 3.4 Theoretical Justification for STAR-MD Architecture
In this subsection, we use the Mori-Zwanzig formalism [mori1965transport, zwanzig1961memory] (See Section˜7 for more details) to provide theoretical justification for two key aspects of our architecture: the necessity of temporal history in coarse-grained modeling, and the specific requirement for joint spatio-temporal attention arising from the removal of explicit pairwise features.
@@
Existing models typically rely on explicit pairwise features to capture spatial structure, but this incurs prohibitive or costs (see Section˜3.2). STAR-MD circumvents this by projecting out pairwise features, a design choice we analyze through our Memory Inflation proposition (Section˜7.2.3, Section˜7). We show that removing explicit spatial correlations “inflates” the memory kernel for the remaining residues, necessitating a significantly richer temporal history to compensate. Crucially, this inflated kernel exhibits non-separable spatio-temporal coupling (Remark˜2), meaning spatial and temporal modes cannot be factorized. This theoretical insight directly motivates our architecture: instead of the interleaved spatial and temporal blocks used in prior work, STAR-MD employs joint spatio-temporal attention to approximate this complex, non-separable memory kernel, balancing physical fidelity with computational scalability.
@@ 4.1 Experimental Setup
Implementations. We follow the training procedure described in Section 3.3 and generate trajectories using the configurations listed in Table 4. We compare STAR-MD with three state-of-the-art trajectory generative models trained on ATLAS: AlphaFolding [cheng20254d], MDGen [jing2024generative_mdgen], and ConfRover [shen2025simultaneous_confrover]. All models parametrize proteins using backbone rigids (translations and rotations) with torsional angles for side-chain atoms. To standardize evaluation, trajectories from variable-stride models (STAR-MD, ConfRover) are generated directly at the required intervals. Trajectories from fixed-stride models (AlphaFolding, MDGen) are first generated at their native resolution and then subsampled. Finally, we include an oracle reference based on MD simulations run independently to represent the target performance. All trajectories are aligned to the starting frame of the reference ATLAS trajectory via C superposition prior to analysis (Section˜10). Specific parameters for each benchmark are detailed in the relevant sections below, with full implementation details in Appendix 12.
@@ 4.5 Ablation Studies on Key Components
To validate our key design choices, we conduct systematic ablations of the main components of STAR-MD. Table 3 summarizes the quantitative results on the benchmark. Section˜14.2 contains full ablation tables for the , , and benchmarks.

### v2

@@ 3.2 Architecture
This architectural choice is key to our model’s scalability. We analyze the per-layer computational complexity of our ST attention () against general ST (i.e., “space-then-time”) architectural paradigms (see Section˜9 for details). Most baseline architectures employ a Pairformer [jumper2021_af2] backbone, incurring a cubic spatial complexity due to expensive triangular attention operations. We consider two representative baseline configurations: (1) Pairformer + Pair Temporal Attention (e.g., ConfRover [shen2025simultaneous_confrover]): This configuration scales as . Here, the pairwise temporal attention alone incurs cost, matching that of STAR-MD, but the additional cubic spatial overhead makes it significantly more expensive. Furthermore, this approach incurs a quadratic memory cost for caching pairwise key-value states during autoregressive generation, which becomes a severe bottleneck for long rollouts. In contrast, STAR-MD only requires memory for caching single-residue features. (2) Pairformer + Single Temporal Attention: This configuration scales as . While the temporal scaling is theoretically superior to STAR-MD’s in the limit , the cubic spatial cost dominates in practical regimes. STAR-MD thus offers a favorable trade-off, enabling efficient modeling of long trajectories for large proteins without the cubic spatial overhead.
@@ 3.4 Theoretical Justification for STAR-MD Architecture
In this subsection, we use the Mori-Zwanzig formalism [mori1965transport, zwanzig1961memory] (See Section˜8 for more details) to provide theoretical justification for two key aspects of our architecture: the necessity of temporal history in coarse-grained modeling, and the specific requirement for joint spatio-temporal attention arising from the removal of explicit pairwise features.
@@
Existing models typically rely on explicit pairwise features to capture spatial structure, but this incurs prohibitive or costs (see Section˜3.2). STAR-MD circumvents this by projecting out pairwise features, a design choice we analyze through our Memory Inflation proposition (Section˜8.2.3, Section˜8). We show that removing explicit spatial correlations “inflates” the memory kernel for the remaining residues, necessitating a significantly richer temporal history to compensate. Crucially, this inflated kernel exhibits non-separable spatio-temporal coupling (Remark˜2), meaning spatial and temporal modes cannot be factorized. This theoretical insight directly motivates our architecture: instead of the interleaved spatial and temporal blocks used in prior work, STAR-MD employs joint spatio-temporal attention to approximate this complex, non-separable memory kernel, balancing physical fidelity with computational scalability.
@@ 4.1 Experimental Setup
Implementations. We follow the training procedure described in Section 3.3 and generate trajectories using the configurations listed in Table 4. We compare STAR-MD with three state-of-the-art trajectory generative models trained on ATLAS: AlphaFolding [cheng20254d], MDGen [jing2024generative_mdgen], and ConfRover [shen2025simultaneous_confrover]. All models parametrize proteins using backbone rigids (translations and rotations) with torsional angles for side-chain atoms. To standardize evaluation, trajectories from variable-stride models (STAR-MD, ConfRover) are generated directly at the required intervals. Trajectories from fixed-stride models (AlphaFolding, MDGen) are first generated at their native resolution and then subsampled. Finally, we include an oracle reference based on MD simulations run independently to represent the target performance. All trajectories are aligned to the starting frame of the reference ATLAS trajectory via C superposition prior to analysis (Section˜11). Specific parameters for each benchmark are detailed in the relevant sections below, with full implementation details in Appendix 13.
@@ 4.5 Ablation Studies on Key Components
To validate our key design choices, we conduct systematic ablations of the main components of STAR-MD. Table 3 summarizes the quantitative results on the benchmark. Section˜15.2 contains full ablation tables for the , , and benchmarks.
@@ 5 Conclusion
6 Author Contributions
Y.S. and Q.G. conceived the study, designed the experimental protocols, and supervised methodology development. N.S., Y.L., and Y.S. developed the spatiotemporal causal diffusion architecture and analyzed model performance. P.L. and Y.S. conceptualized the contextual noise perturbation, which N.S. implemented. N.S., P.L. and R.B. conducted the theoretical verification of the Mori-Zwanzig formalism and validated the mathematical frameworks. All authors participated in the manuscript preparation and critical review for publication.